OpenStudy (ammarah):

Solve the equation by finding square roots. 3(x-3)^+2=26

OpenStudy (ammarah):

@Easyaspi314

ganeshie8 (ganeshie8):

start by subtracting 2 both sides

OpenStudy (ammarah):

ok then what

ganeshie8 (ganeshie8):

then few more steps, wat do u get when u subtract ?

OpenStudy (ammarah):

3(x-3)^=24

ganeshie8 (ganeshie8):

then divide 3 both sides

OpenStudy (ammarah):

wait the swaure root of 8 is 4 rad 2 ??

ganeshie8 (ganeshie8):

sqrt(8) = sqrt(4*2) = 2sqrt(2)

OpenStudy (ammarah):

i dont get this one: 5x^+13x=2x^+4

ganeshie8 (ganeshie8):

is this a different question.... ? are we done wid first question ?

OpenStudy (ammarah):

ya

ganeshie8 (ganeshie8):

may i knw wat did u get as solution ?

ganeshie8 (ganeshie8):

cuz there are two solutions, and i think u seem to be missing one

OpenStudy (ammarah):

? for the first question i got 3 plus or minus 2 rad 2

ganeshie8 (ganeshie8):

okay we're good :)

OpenStudy (ammarah):

can u please help me with this next one i have to go

ganeshie8 (ganeshie8):

5x^2+13x=2x^2+4

ganeshie8 (ganeshie8):

subtract 2x^2 both sides

ganeshie8 (ganeshie8):

then complete the square

ganeshie8 (ganeshie8):

give it a try

OpenStudy (ammarah):

wait so it would be 13/2 and then u take the suare root???

ganeshie8 (ganeshie8):

not that simple

ganeshie8 (ganeshie8):

before we begin solving, u sure ur equation is really like this ? 5x^2+13x=2x^2+4

OpenStudy (ammarah):

wait no its 13 not 13x

ganeshie8 (ganeshie8):

5x^2+13=2x^2+4 like this ?

OpenStudy (ammarah):

yes

ganeshie8 (ganeshie8):

5x^2+13=2x^2+4 subtract 2x^2 both sides 3x^2 + 13 = 4 subtract 13 both sides 3x^2 = -9 divide 3 both sides x^2 = -3

ganeshie8 (ganeshie8):

then what ?